a metetologist kept track of number of days it snowed in january thresults are
4 5. 12 9 10 6 5 8 15 9
find the 95% confidence interval of mean of days it snowe? round to the nearest whole number?
Answer #1

Solution :

MEAN OF OBSERVATIONS = SUM OF OBSERVATIONS /NUMBER OF OBSERVATIONS

MEAN OF OBSERVATIONS = 83/10 = 8.3

STANDARD DEVIATION = 3.28

CONFIDENCE LEVEL 95%

M = 8.3
t = 1.96
sM= √(3.282/10) = 1.04

MEAN = M ± Z(sM)
MEAN = 8.3 ± 1.96*1.04
MEAN = 8.3 ± 2.033

M = 8.3, 95% CI [6.267, 10.333].

You can be 95% confident that the population mean (μ) falls between 6.267 and 10.333.